Dear,

Here is a summary of CREAM-IV flight preparation status on the 14th of November at McMurdo.

  • For most of day, TCD/CD data were taken with different threshold values for the calibration of the threshold values.
  • During the calibration run, PMT #1 anode problem with low signal found and fixed by replacing the signal cable between PMT and Concentrator box (PXL).
  • Chercam reset coding to flush data in the Chercam sparsification board was prepared by Prashant and tested working fine by removing event # cable. Before this, the Chercam data was not flushed so that event number mismatching happened continuously until the Chercam event number is matching with others. Since there was no waiting time but just throwing away those events with event number mismatching, this was not found to be critical until we tried TCD time out debugging. After adding the reset code, event mismatch by Chercam reduced.
  • Rich (NSF) has put two thermometers with data logger in the building for monitoring temperatures.
  • Fritz worked on fit check of the foam walls.
  • Nothing special in the weather briefing. The wind map at 5 mbar altitude which is now inside the CSBF internal web page is going to be available to the public soon.

UMD